# Solar Neutrinos and the Strongest Oscillation Constraints on Scalar NSI

### v1

@@ 1 Introduction
While NSI’s with different Lorentz structures have been examined in scattering for some time, only recently has there become a growing interesting in scalar NSI (sNSI) in neutrino oscillation environments Ge:2018uhz ; Babu:2019iml ; Smirnov:2019cae ; Arguelles:2019xgp ; Venzor:2020ova ; Medhi:2021wxj ; Medhi:2022qmu ; Sarkar:2022ujy ; Dutta:2022fdt ; Denton:2022pxt ; Medhi:2023ebi ; Singha:2023set ; Sarker:2023qzp ; Sarker:2024ytu ; Dutta:2024hqq ; ESSnuSB:2023lbg ; Gupta:2023wct . Notably Babu:2019iml investigates the interplay of oscillation and non-oscillation constraints on sNSI under a variety of assumptions about the mediator mass, the nature of neutrinos, and whether the interaction is with electrons or nucleons. There are key regions of parameter space where constraints from oscillations, particularly solar neutrinos, dominate. The solar constraints were derived in Ge:2018uhz by looking only at solar neutrino data from Borexino Borexino:2017rsf . Additional solar neutrino data sets exist, however, some of which are considerably more precise. In addition, nearly all existing studies assumed the lightest neutrino mass, $m_{1}$ in the normal ordering (NO) and $m_{3}$ in the inverted ordering (IO) to be exactly massless [^4] Ref. Medhi:2023ebi investigated the impact of neutrino masses in sNSI at future long-baseline accelerator neutrino experiments..
@@ 2 Scalar NSI
where $a=2\sqrt{2}G_{F}N_{e}E$, $G_{F}$ is Fermi’s constant, and $N_{e}$ is the electron number density, while $\tilde{M}$ is defined as
@@ 2.1 Scalar NSI in the Sun
We want now to understand which is the effects of the scalar NSI on the solar neutrino survival probabilities. Since, as already pointed out, the scalar NSI parameters depends on the matter density, neutrino produced at different radii in the Sun are affected differently by such interactions. We assume that $\phi$ couples equally to up and down quarks and not to electrons, thus the $\sum_{f}N_{f}y_{f}$ sum is proportional to the density. A different flavor structure for the couplings to matter fermions will change the picture somewhat because the neutron fraction evolves with radius in the Sun and different neutrino populations are produced dominantly at different radii, although this effect is not large. Thus, we further parametrize the new physics parameters in the following way:
@@
Figure 1: Survival probability for solar neutrinos as a function of the lightest neutrino mass $m_{1}$ and $\eta_{ee}^{\odot}$ (left), $\eta_{\mu\mu}^{\odot}$ (middle) and $\eta_{\tau\tau}^{\odot}$ (right). In the top panels we show the probability for pp neutrinos while in the bottom panels the probability for ${}^{8}B$ neutrinos. The dashed curves represents the region of the parameter space allowed by current measurements.
Figure 2: Survival probability for solar neutrinos as a function of $\eta_{e\mu}^{\odot}$ (left) $\eta_{e\tau}^{\odot}$ (middle) and $\eta_{\mu\tau}^{\odot}$ (right) and their phases for $m_{1}=0$ eV. In the top panels we show the probability for pp neutrinos while in the bottom ones the probability for ${}^{8}B$ neutrinos.
@@ 3 Solar neutrinos
  • ${}^{7}Be$ neutrinos: neutrinos produced by the Beryllium 7 electron capture. These neutrinos have a fixed energy of 0.826 MeV
@@
  • ${}^{8}B$ neutrinos: in this case neutrinos are produced by the $\beta$-decays of boron nuclei. Their flux is four order of magnitude smaller than the pp one but the mean neutrino energy is around 8.1 MeV.
These four kind of solar neutrinos are particularly interesting because they are affected in different ways by the matter potential inside the Sun due to their range of energies. Indeed, while for the less energetic pp neutrinos the Sun does not affect their oscillations in the conventional picture, the ${}^{8}B$ neutrino undergo the MSW transition due to the presence of matter in the usual picture. On the other hand, the intermediate pep neutrinos have an energy which allow us to probe the energy region in which the transition occurs. In the adiabatic approximation, characterized by slowly changing matter density, it is possible to compute the oscillation disappearance probability for a terrestrial experiment. In the two flavor approximation, averaging out the oscillation occurring during the long distance between the Sun and the Earth and neglecting the so-called “jump probability” Parke:1986jy ; Haxton:1986dm ; Petcov:1987xd , we have:
@@
that is what we observe for more energetic neutrinos like the ${}^{8}B$ neutrinos.
@@
where $R$ is the Sun radius and $\phi(r)$ is a normalized function describing where neutrinos are produced in the Sun. In our analysis we use the BS05(OP) model Bahcall:2004pz . The survival probability in eq. 13 is then slightly modified to:
@@
$P_{ee}=\cos^{4}\theta_{13}\frac{1}{2}P_{ee}^{2\nu}+\sin^{4}\theta_{13}\,,$ (20)
where $P_{ee}^{2\nu}$ is the two-flavor probability in eq. 19 and $\cos^{2}\theta_{13}\tilde{A}$ as the relevant three-flavor matter potential.
@@
where $U_{ei}^{s}$ are the mixing matrix elements in matter and $U_{ei}$ are those in vacuum. In table 2 we report the values of the survival probabilities for the four solar neutrino components computed using the 2-flavor approximation, the 2-flavor formula with the correction and the 3-flavor full probability. The values of the oscillation parameters are taken from table 1. In Fig. 4 we compute for a fixed neutrino energy corresponding to the four solar neutrino components, the value of the three mixing matrix elements $|U_{e1}|^{2}$, $|U_{e2}|^{2}$, and $|U_{e3}|^{2}$ as a function of the neutrino production point inside the Sun. We also show with dashed lines the values of such matrix elements in vacuum and we superimpose the neutrino production functions $\phi(r)$ in arbitrary units. It is clear that for less energetic neutrinos (pp neutrinos), the matter potential mildly modifies the vacuum values of the mixing matrix elements. On the other hand, for pep neutrinos, at an energy of around 1.5 MeV, the transition between the two matrix elements $|U_{e1}|^{2}$ and $|U_{e2}|^{2}$ occurs at a radius which is around the production points of pep neutrinos. Finally, for the ${}^{8}B$ neutrinos at a higher energy (8.1 MeV) the matter transition has completely occurred: $|U_{e1}|^{2}$ is close to one, while $|U_{e2}|^{2}$ is almost 0. It is worth to mention that, as expected, the 3-flavors correction on $|U_{e3}|^{2}$ remains $\sim 2\%$, largely independent of the radius and density, as expected.
@@
Table 2: Expected survival probabilities computed for the three solar neutrino components using eq. 19 (left column), 20 (middle column) and 21 (right column). Values for the oscillation parameters are taken from table 1.
@@
${}^{8}B$ 0.31 0.30 0.30
Figure 4: Values of the mixing matrix element $|U_{ei}|^{2}$ in matter as a function of the neutrino production point for different neutrino energies, corresponding to pp neutrinos (top left panel), ${}^{7}Be$ neutrinos (top right panel), pep neutrinos (bottom left panel) and ${}^{8}B$ neutrinos (bottom right panel). We also show with dashed lines the values of the mixing matrix elements in vacuum and the production point functions for the three neutrino components in arbitrary units (yellow region).
@@ 3.1 Solar neutrino oscillation in presence of scalar NSI
In this section we will discuss the effect of scalar NSI on the solar neutrino oscillation. In general, the modification to the oscillation probabilities given by scalar NSI are very complicated to evaluate. Indeed, scalar NSI parameters can drastically alter the neutrino flavor transition; for instance, in presence of matter, a combination of $\eta_{\alpha\beta}$ might mimic the effect of the neutrino mass splittings and produce the observed oscillatory pattern even if the neutrino mass states are degenerate. In Fig. 1 we show the survival probabilities $P_{ee}$ in the $\eta_{\alpha\alpha}^{\odot}$ vs $m_{1}$ plane, where $m_{1}$ is the lightest neutrino mass (NO is assumed here). We show them for pp and $8^{B}$ neutrinos. Furthermore, in Figs. 2 and 3 we show the survival probabilities in the planes $\eta_{\alpha\beta}^{\odot}$ vs. $\phi_{\alpha\beta}$ obtained for $m_{1}=0$ eV and $m_{1}=0.01$ eV, respectively. In all three figures, we show with dashed lines the parameters spaces for which we obtain survival probabilities compatible with the experimental ones BOREXINO:2018ohr .
Under some approximations, we can try to understand analitically the effect of the scalar NSI on the effective mass splitting and mixing angle for the solar neutrinos. Firstly, for sake of simplicity, in the following discussion we will take into account only two neutrino flavors. The main disadvantage of this approach is that not all the 6 scalar NSI parameters can be discussed. Conversely, the analytical results in this case are more readable and will allow to draw some conclusions on what to expect from our simulations. The second approximation we want to adopt here is to expand our results in around $\eta=0$ and only take the leading correction to the observables. This again does not allow to catch all the possible effects of scalar NSI, which in principle can also be large. However, in the context of solar neutrinos, we can expect that the least energetic pp-neutrinos are not affected neither by usual matter effects nor by the scalar NSI with solar matter. Thus, the ”vacuum-like” measurement of pp-neutrinos will not allow for large values of $\eta$ and our expansion can be considered valid. Finally, with the rescaling adopted in eq. 10, the scalar NSI parameters could reach values greater than 1; this in principle breaks our series expansion. Nonetheless, the choice of the rescaling density is arbitrary and, again, we can still draw our conclusions following our approach.
@@
where $\tilde{A}$ has been defined in eq. 15 and $m_{1}$, $m_{2}$ are the two neutrino absolute masses (with $\Delta m^{2}=m_{2}^{2}-m_{1}^{2}$). In the simple limit $m_{1}\to 0$, this correction always has the same sign of $\eta_{1}$. Since we do observe the matter transition in solar data, with a depletion of the oscillation probability for high-energy ${}^{8}B$ neutrinos, the effective mass splitting cannot be too large. This set a stringent upper limit on $\eta_{1}$. However, if $\eta_{1}<0$, data can still be fitted with a larger $\Delta m^{2}$. For this reason, the lower limit on $\eta_{1}$ set by solar data is less stringent than the upper one for $m_{1}\to 0$. In the case of a non-vanishing $m_{1}$, on the other hand, for high energetic solar neutrinos, $\tilde{A}>1$ and the correction has the same $\eta_{1}$ sign but it is larger, reducing the allowed values for the scalar NSI parameters. Let us now consider the effective mixing angle in presence of $\eta_{1}$. In this case we obtain the correction
@@
This correction, for $m_{1}\to 0$, has again the opposite $\eta_{1}$ sign. Given that $\cos 2\theta_{12}^{s}\to-1$ for ${}^{8}B$ neutrinos, we can observe that the disappearance probability in eq. 19 can be fitted with smaller values of $\theta_{12}$ if $\eta_{1}>0$. Larger values of the mixing angle cannot easily accommodate the data since they would significantly increase $\Delta m^{2}_{21,\rm{eff}}$ pushing the matter transition towards higher energies. An increase of $m_{1}$ also increases the value of the effective mixing angle.
Let us now consider the case in which $\eta_{2}\neq 0$. We obtain for the mass splitting and the mixing angle the following corrections
@@
The correction to the mass splitting has again the same sign of the parameter $\eta_{3}$. This again prefers negative values for $\eta_{3}$ (which correspond to $\phi=\pi$ in the case of a complex parameter). On the other hand, the correction to the mixing angle has an opposite sign with respect to $\eta_{3}$ for neutrinos which undergo matter transition. In this case, as in the $\eta_{1}$ case, a fit with smaller $\theta_{12}$ values is viable if $\eta_{3}>0$.
@@ 4 Solar data analysis in presence of diagonal scalar NSI parameters
We now want to understand which values of the scalar NSI are allowed by solar neutrino measurements. Starting from the observed values for the electron neutrino disappearance probabilities by Borexino BOREXINO:2018ohr
@@
and adding the spectral information on the ${}^{8}B$ neutrinos by SNO SNO:2011hxd  [^7] Additional solar neutrino information from the dedicated pp experiments SAGE, GALEX, and GNO GALLEX:1998kcz ; SAGE:1999nng ; Gavrin:2001sz ; GNO:2000avz and 8B measurements Super-Kamiokande Super-Kamiokande:2016yck will add some additional information, but will not significantly enhance the constraint., we then compute the $\chi^{2}$ function defined as
@@
where $p$ are all the neutrino components already mentioned, $P_{obs}^{p}$ are the observed probabilities with their uncertainties $(\sigma^{p}_{obs})^{2}$ in eq. 27 and $P(\theta_{12},\Delta m_{21}^{2},\eta,\phi,m)$ are the probabilities computed in presence of scalar NSI fixing the $\theta_{13}$ value to its best fit Esteban:2020cvm using the exact probabilities from eq. 21.
Figure 5: 90% C.L. contours in the $\eta_{\alpha\alpha}^{\odot}$ vs $m_{1}$ planes for the three diagonal scalar NSI parameters using solar neutrino data after minimizing over the relevant oscillation parameters. The stars depict the best fit points.
In Fig. 5 we show the 90% CL allowed regions in the $\eta^{\odot}_{\alpha\alpha}-m_{1}$ planes marginalizing over the standard oscillation parameters [^8] We consider for $\Delta m_{21}^{2}$ only values in the range $[10^{-5}-10^{-4}]$ $eV^{2}$. It is clear that if the lightest neutrino mass is large ($m_{1}>10^{-2}$ eV), the scalar NSI parameters are strongly constrained. Indeed, for those values the new physics effects would be too large even for small $\eta_{\alpha\alpha}^{\odot}$ to reproduce the experimental measurements. However, if the lightest neutrino mass is small enough, data could be still reproduced if $\eta_{ee}^{\odot}\in[-2,0.5]$, $\eta_{\mu\mu,\tau\tau}^{\odot}\in[-3,1]$. In the valid approximation $\eta_{1}\to\eta_{ee}^{\odot}$ and $\eta_{2}\to\eta_{ee}^{\odot}$ it is possible to see that, as expected from the results shown in Sec. 3.1, negative values of the $\eta_{\alpha\alpha}^{\odot}$ are preferred since, in the case of large positive values, the matter transition would not occur for ${}^{8}B$ neutrinos.

### v2

@@ 1 Introduction
While NSI’s with different Lorentz structures have been examined in scattering for some time, only recently has there become a growing interesting in scalar NSI (sNSI) in neutrino oscillation environments Ge:2018uhz ; Babu:2019iml ; Smirnov:2019cae ; Arguelles:2019xgp ; Suliga:2020lir ; Venzor:2020ova ; Medhi:2021wxj ; Medhi:2022qmu ; Sarkar:2022ujy ; Dutta:2022fdt ; Denton:2022pxt ; Medhi:2023ebi ; Singha:2023set ; Sarker:2023qzp ; Sarker:2024ytu ; Dutta:2024hqq ; ESSnuSB:2023lbg ; Gupta:2023wct ; Liao:2015rma . Ref. Suliga:2020lir investigated new scalar (and vector) interactions in the Sun, but not in the context of oscillations. Notably Babu:2019iml investigates the interplay of oscillation and non-oscillation constraints on sNSI under a variety of assumptions about the mediator mass, the nature of neutrinos, and whether the interaction is with electrons or nucleons. There are key regions of parameter space where constraints from oscillations, particularly solar neutrinos, dominate. The solar constraints were derived in Ge:2018uhz by looking only at solar neutrino data from Borexino Borexino:2017rsf . Additional solar neutrino data sets exist, however, some of which are considerably more precise. In addition, nearly all existing studies assumed the lightest neutrino mass, $m_{1}$ in the normal ordering (NO) and $m_{3}$ in the inverted ordering (IO) to be exactly massless [^4] Ref. Medhi:2023ebi investigated the impact of neutrino masses in sNSI at future long-baseline accelerator neutrino experiments..
@@ 2 Scalar NSI
where
$a=2\sqrt{2}G_{F}N_{e}E\,,$ (9)
$G_{F}$ is Fermi’s constant, and $N_{e}$ is the electron number density, while $\tilde{M}$ is defined as
@@ 2.1 Scalar NSI in the Sun
We want now to understand what is the effects of the scalar NSI on the solar neutrino survival probabilities. Since, as already pointed out, the scalar NSI parameters depends on the matter density, neutrino produced at different radii in the Sun are affected differently by such interactions. We assume that $\phi$ couples equally to up and down quarks and not to electrons, thus the $\sum_{f}N_{f}y_{f}$ sum is proportional to the density. A different flavor structure for the couplings to matter fermions will change the picture somewhat because the neutron fraction evolves with radius in the Sun and different neutrino populations are produced dominantly at different radii, although this effect is not large. Thus, we further parametrize the new physics parameters in the following way:
@@
Figure 1: Survival probability for solar neutrinos as a function of the lightest neutrino mass $m_{1}$ and $\eta_{ee}^{\odot}$ (left), $\eta_{\mu\mu}^{\odot}$ (middle) and $\eta_{\tau\tau}^{\odot}$ (right). In the top panels we show the probability for pp neutrinos while in the bottom panels the probability for 8B neutrinos. The dashed curves represents the region of the parameter space allowed by current measurements.
Figure 2: Survival probability for solar neutrinos as a function of $\eta_{e\mu}^{\odot}$ (left) $\eta_{e\tau}^{\odot}$ (middle) and $\eta_{\mu\tau}^{\odot}$ (right) and their phases for $m_{1}=0$ eV. In the top panels we show the probability for pp neutrinos while in the bottom ones the probability for 8B neutrinos.
@@ 3 Solar neutrinos
  • 7Be neutrinos: neutrinos produced by the Beryllium 7 electron capture. These neutrinos have a fixed energy of 0.826 MeV
@@
  • 8B neutrinos: in this case neutrinos are produced by the $\beta$-decays of boron nuclei. Their flux is four order of magnitude smaller than the pp one but the mean neutrino energy is around 8.1 MeV.
These four kind of solar neutrinos are particularly interesting because they are affected in different ways by the matter potential inside the Sun due to their range of energies. Indeed, while for the less energetic pp neutrinos the Sun does not affect their oscillations in the conventional picture, the 8B neutrino undergo the MSW transition due to the presence of matter in the usual picture. On the other hand, the intermediate pep neutrinos have an energy which allow us to probe the energy region in which the transition occurs. In the adiabatic approximation, characterized by slowly changing matter density, it is possible to compute the oscillation disappearance probability for a terrestrial experiment. In the two flavor approximation, averaging out the oscillation occurring during the long distance between the Sun and the Earth and neglecting the so-called “jump probability” Parke:1986jy ; Haxton:1986dm ; Petcov:1987xd , we have:
@@
that is what we observe for more energetic neutrinos like the 8B neutrinos.
@@
where $R$ is the Sun radius and $\phi(r)$ is a normalized function describing where neutrinos are produced in the Sun. The radial dependence in $\theta_{12}^{s}(r)$ comes from the fact that $\theta_{12}^{s}$ depends on the matter effect (see eq. 15) which in turn depends on the electron density (see eqs. 16 and 9), and the electron density depends on the radius of the Sun as given by the Standard Solar Model. In our analysis we use the BS05(OP) model Bahcall:2004pz . The survival probability in eq. 14 is then slightly modified to:
@@
$P_{ee}=\cos^{4}\theta_{13}P_{ee}^{2\nu}+\sin^{4}\theta_{13}\,,$ (21)
where $P_{ee}^{2\nu}$ is the two-flavor probability in eq. 20 and $\cos^{2}\theta_{13}\tilde{A}$ is the relevant three-flavor matter potential. See e.g. Lim:1987yd ; Akhmedov:1999uz .
@@
where $U_{ei}^{s}$ are the mixing matrix elements in matter and $U_{ei}$ are those in vacuum. In table 2 we report the values of the survival probabilities for the four solar neutrino components computed using the 2-flavor approximation, the 2-flavor formula with the correction and the 3-flavor full probability. The values of the oscillation parameters are taken from table 1. In Fig. 4 we compute for a fixed neutrino energy corresponding to the four solar neutrino components, the value of the three mixing matrix elements $|U_{e1}|^{2}$, $|U_{e2}|^{2}$, and $|U_{e3}|^{2}$ as a function of the neutrino production point inside the Sun. We also show with dashed lines the values of such matrix elements in vacuum and we superimpose the neutrino production functions $\phi(r)$ in arbitrary units. It is clear that for less energetic neutrinos (pp neutrinos), the matter potential mildly modifies the vacuum values of the mixing matrix elements. On the other hand, for pep neutrinos, at an energy of around 1.5 MeV, the transition between the two matrix elements $|U_{e1}|^{2}$ and $|U_{e2}|^{2}$ occurs at a radius which is around the production points of pep neutrinos. Finally, for the 8B neutrinos at a higher energy (8.1 MeV) the matter transition has completely occurred: $|U_{e1}|^{2}$ is close to one, while $|U_{e2}|^{2}$ is almost 0. It is worth to mention that, as expected, the 3-flavors correction on $|U_{e3}|^{2}$ remains $\sim 2\%$, largely independent of the radius and density, as expected.
@@
Table 2: Expected survival probabilities computed for the three solar neutrino components using eq. 20 (left column), 21 (middle column) and 22 (right column). Values for the oscillation parameters are taken from table 1.
@@
8B 0.31 0.30 0.30
Figure 4: Values of the mixing matrix element $|U_{ei}|^{2}$ in matter as a function of the neutrino production point for different neutrino energies, corresponding to pp neutrinos (top left panel), 7Be neutrinos (top right panel), pep neutrinos (bottom left panel) and 8B neutrinos (bottom right panel). We also show with dashed lines the values of the mixing matrix elements in vacuum and the production point functions for the three neutrino components in arbitrary units (yellow region).
@@ 3.1 Solar neutrino oscillation in presence of scalar NSI
In this section we will discuss the effect of scalar NSI on the solar neutrino oscillation. In general, the modification to the oscillation probabilities given by scalar NSI are very complicated to evaluate. Indeed, scalar NSI parameters can drastically alter the neutrino flavor transition; for instance, in presence of matter, a combination of $\eta_{\alpha\beta}$ might mimic the effect of the neutrino mass splittings and produce the observed oscillatory pattern even if the neutrino mass states are degenerate. In Fig. 1 we show the survival probabilities $P_{ee}$ in the $\eta_{\alpha\alpha}^{\odot}$ vs $m_{1}$ plane, where $m_{1}$ is the lightest neutrino mass (NO is assumed here). We show them for pp and 8B neutrinos. Furthermore, in Figs. 2 and 3 we show the survival probabilities in the planes $\eta_{\alpha\beta}^{\odot}$ vs. $\phi_{\alpha\beta}$ obtained for $m_{1}=0$ eV and $m_{1}=0.01$ eV, respectively. In all three figures, we show with dashed lines the parameters spaces for which we obtain survival probabilities compatible with the experimental ones BOREXINO:2018ohr .
To understand the physics of sNSI in the Sun, we develop some approximations; the numerical work below is carried out using the full calculations. Under some approximations, we can try to understand analytically the effect of the scalar NSI on the effective mass splitting and mixing angle for the solar neutrinos. First, for sake of simplicity, in the following discussion we will take into account only two neutrino flavors. The analytical results in this case are more readable and will allow to draw some conclusions on what to expect from our simulations. The main disadvantage of this approach, however, is that not all the 6 scalar NSI parameters can be discussed. The second approximation we adopt in this subsection is to expand our results in around $\eta=0$ and only take the leading correction to the observables. This again does not allow to catch all the possible effects of scalar NSI, which in principle can be large. However, in the context of solar neutrinos, we can expect that the least energetic pp-neutrinos are not affected neither by usual matter effects nor by the scalar NSI with solar matter, in part due to the lower energies and in part because pp neutrinos come from larger radii, and thus lower densities, than e.g. 8B neutrinos. Thus, large values of the $\eta$ parameters will not be too relevant for the measurement of pp-neutrinos and our expansion is quite reliable in some useful regimes. Finally, with the rescaling adopted in eq. 11, the scalar NSI parameters could reach values greater than 1; this in principle breaks our series expansion. Nonetheless, the choice of the rescaling density is arbitrary and, again, we can still draw our conclusions following our approach.
@@
where $\tilde{A}$ has been defined in eq. 16 and $m_{1}$, $m_{2}$ are the two neutrino absolute masses (with $\Delta m^{2}=m_{2}^{2}-m_{1}^{2}$). In the simple limit $m_{1}\to 0$, this correction always has the same sign as $\eta_{1}$. Since we do observe the matter transition in solar data, with a depletion of the oscillation probability for high-energy 8B neutrinos, the effective mass splitting cannot be too large. This set a stringent upper limit on $\eta_{1}$. However, if $\eta_{1}<0$, data can still be fitted with a larger $\Delta m^{2}$. For this reason, the lower limit on $\eta_{1}$ set by solar data is less stringent than the upper one for $m_{1}\to 0$. In the case of a non-vanishing $m_{1}$, on the other hand, for high energetic solar neutrinos, $\tilde{A}>1$ and the correction has the same sign as $\eta_{1}$, but it is larger, reducing the allowed values for the scalar NSI parameters. Let us now consider the effective mixing angle in presence of $\eta_{1}$. In this case we obtain the correction
@@
This correction, for $m_{1}\to 0$, has again the opposite sign as $\eta_{1}$. Given that $\cos 2\theta_{12}^{s}\to-1$ for 8B neutrinos, we can observe that the disappearance probability in eq. 20 can be fitted with smaller values of $\theta_{12}$ if $\eta_{1}>0$. Larger values of the mixing angle cannot easily accommodate the data since they would significantly increase $\Delta m^{2}_{21,\rm{eff}}$ pushing the matter transition towards higher energies. An increase of $m_{1}$ also increases the value of the effective mixing angle.
Let us now consider the case in which $\eta_{2}\neq 0$. We obtain the following corrections for the mass splitting and the mixing angle
@@
The correction to the mass splitting has again the same sign of the parameter $\eta_{3}$. This again prefers negative values for $\eta_{3}$ (which correspond to $\phi=\pi$ in the case of a complex parameter). On the other hand, the correction to the mixing angle has the opposite sign with respect to $\eta_{3}$ for neutrinos which undergo matter transition. In this case, as in the $\eta_{1}$ case, a fit with smaller $\theta_{12}$ values is viable if $\eta_{3}>0$.
@@ 4 Solar data analysis in presence of diagonal scalar NSI parameters
We now want to understand which values of the scalar NSI are allowed by solar neutrino measurements using the full-three flavor picture and for arbitrary $\eta$ values (that is, we are not following either of the simplifying approximations from the previous section). Starting from the observed values for the electron neutrino disappearance probabilities by Borexino BOREXINO:2018ohr
@@
and adding the spectral information on the 8B neutrinos by SNO SNO:2011hxd  [^7] Additional solar neutrino information from the dedicated pp experiments SAGE, GALEX, and GNO GALLEX:1998kcz ; SAGE:1999nng ; Gavrin:2001sz ; GNO:2000avz and 8B measurements Super-Kamiokande Super-Kamiokande:2016yck will add some additional information, but will not significantly enhance the constraint. including uncertainties, we then compute the $\chi^{2}$ function defined as
@@
where $p$ are all the neutrino components already mentioned, $P_{obs}^{p}$ are the observed probabilities with their uncertainties $(\sigma^{p}_{obs})^{2}$ in eq. 28 and $P(\theta_{12},\Delta m_{21}^{2},\eta,\phi,m)$ are the probabilities computed in presence of scalar NSI using the exact probabilities from eq. 22 fixing $\theta_{13}$ and the other oscillation parameters in the atmospheric sector ($\theta_{23}$, $\Delta m^{2}_{31}$) fixed to their best fits (see Tab. 2 Esteban:2020cvm ) while setting CP violation to 0. We have checked that in the solar oscillation regime the effect of the atmospheric parameters, of $\delta_{\rm CP}$, and of the reactor mixing angle is negligible even in presence of scalar NSI.
Figure 5: 90% CL contours in the $\eta_{\alpha\alpha}^{\odot}$ vs $m_{1}$ planes for the three diagonal scalar NSI parameters using solar neutrino data after minimizing over the relevant oscillation parameters. The stars depict the best fit points.
In Fig. 5 we show the 90% CL allowed regions at 2 degrees of freedom in the $\eta^{\odot}_{\alpha\alpha}-m_{1}$ planes marginalizing [^8] For $\Delta m_{21}^{2}$ we consider only values in the range $[10^{-5}-10^{-4}]$ $eV^{2}$. over the standard solar oscillation parameters: $\theta_{12}$ and $\Delta m^{2}_{21}$ while keeping the other sNSI parameters fixed to zero. It is clear that if the lightest neutrino mass is large ($m_{1}>10^{-2}$ eV), the scalar NSI parameters are strongly constrained. Indeed, for those values the new physics effects would be too large even for small $\eta_{\alpha\alpha}^{\odot}$ to reproduce the experimental measurements. However, if the lightest neutrino mass is small enough, data could be still reproduced if $\eta_{ee}^{\odot}\in[-2,0.5]$, $\eta_{\mu\mu,\tau\tau}^{\odot}\in[-3,1]$. In the valid approximation $\eta_{1}\to\eta_{ee}^{\odot}$ and $\eta_{2}\to\eta_{ee}^{\odot}$ it is possible to see that, as expected from the results shown in Sec. 3.1, negative values of the $\eta_{\alpha\alpha}^{\odot}$ are preferred since, in the case of large positive values, the matter transition would not occur for 8B neutrinos.